OpenStudy (anonymous):

f(x) = x2 - 8x + 5. Find f(-1).

OpenStudy (amistre64):

what part of this is giving you a problem?

OpenStudy (anonymous):

(-1)^2-8(-1)+5 =1+8+5 =14